### v1

# CoAM: Corpus of All-Type Multiword Expressions

Abstract
Multiword expressions (MWEs) refer to idiomatic sequences of multiple words. MWE identification, i.e., detecting MWEs in text, can play a key role in downstream tasks such as machine translation. Existing datasets for MWE identification are inconsistently annotated, limited to a single type of MWE, or limited in size. To enable reliable and comprehensive evaluation, we created CoAM: Corpus of All-Type Multiword Expressions, a dataset of 1.3K sentences constructed through a multi-step process to enhance data quality consisting of human annotation, human review, and automated consistency checking. MWEs in CoAM are tagged with MWE types, such as Noun and Verb, to enable fine-grained error analysis. Annotations for CoAM were collected using a new interface created with our interface generator, which allows easy and flexible annotation of MWEs in any form, including discontinuous ones.111 The CoAM dataset and source code of the interface generator will be publicly accessible upon acceptance. Through experiments using CoAM, we find that a fine-tuned large language model outperforms the current state-of-the-art approach for MWE identification. Furthermore, analysis using our MWE type tagged data reveals that Verb MWEs are easier than Noun MWEs to identify across approaches.
CoAM: Corpus of All-Type Multiword Expressions
Yusuke Ide1  Joshua Tanner2  Adam Nohejl1  Jacob Hoffman2  Justin Vasselli1 Hidetaka Kamigaito1  Taro Watanabe1 1Nara Institute of Science and Technology  2Resolve Research {ide.yusuke.ja6, nohejl.adam.mt3, vasselli.justin_ray.vk4, kamigaito.h, taro} @is.naist.jp, {josh, jake}@rslv.dev
1 Introduction
Vocabulary plays a critical role in the comprehension of natural language. Vocabulary is often thought of as consisting of single words, but multiword expressions (MWEs) also form an important part (Jackendoff, 1995). MWEs are defined as idiomatic sequences of multiple words Baldwin and Kim (2010). We focus specifically on those whose meaning or grammatical structure cannot be derived from their constituent words, as they can impede text analysis or comprehension. In other words, we exclude transparent collocations. For example, the bold sequences below constitute MWEs.
  1. (a)
    ACL stands for Association for Computational Linguistics.
  2. (b)
    He has been under the weather lately.
MWE identification (MWEI), the process of automatically tagging MWEs in text, has been the object of much research (Constant et al., 2017). MWEI can play a key role in a wide range of downstream tasks, including, but not limited to, machine translation Briakou et al. (2024) and lexical complexity assessment Kochmar et al. (2020). However, as described in Section 2, existing datasets for MWE identification are inconsistently annotated, limited to a single type of MWE, or limited in size. This hinders reliable and comprehensive evaluations of MWEI systems.
In this paper, we introduce the Corpus of All-type Multiword expressions (CoAM), a dataset of 1.3K sentences for MWEI. “Types” refer to MWE categories assigned based on the part of speech of the MWE, with “all-type” signifying the inclusion of all such categories of MWE. To ensure annotation quality, we assigned two annotators and a reviewer to each sentence and checked all annotations for consistency. Additionally, MWEs in the CoAM test set are tagged with their types, such as Noun and Verb, facilitating fine-grained error analysis of MWEI systems. This enables us to address questions such as: “Are verbal MWEs—the focus of the long-running PARSEME project Savary et al. (2017)—more difficult to identify compared to other categories of MWEs?”
Annotations for CoAM were collected using a checkbox-based annotation interface created using our new interface generator, CAIGen: Checkbox-based Annotation Interface Generator. It allows easy and flexible annotation of MWEs in any form, including discontinuous MWEs such as pick … up in pick me up at the station, which are common in English and have historically been a challenge for MWEI systems Rohanian et al. (2019).
Using CoAM, we evaluate two distinct MWEI approaches. The first approach, MWEasWSD (MaW, Tanner and Hoffman, 2023), combines a rule-based pipeline and a trainable bi-encoder model, achieving state-of-the-art performance on the DiMSUM dataset Schneider et al. (2016). The second is LLM fine-tuning for MWEI, inspired by the effectiveness of similar approaches in tasks such as named entity recognition (NER, Zhou et al., 2024). We use CoAM to train and evaluate LLMs from the Llama Dubey et al. (2024) and Qwen Qwen Team (2024) model families.
The results reveal that a fine-tuned Qwen model with 72B parameters greatly outperforms MaW, suggesting the effectiveness of LLM fine-tuning. On the other hand, all approaches suffer from low recall (e.g., 50.7% for Qwen-72B). Further analysis shows that fine-tuned LLMs struggle more to detect Noun and Clause MWEs than to detect Verb MWEs. MWEs not contained in WordNet Miller (1995), e.g., real estate, were found particularly difficult to identify, presumably because they are less widely recognized as MWEs.
2 Related Work
Datasets
Previous studies have presented several datasets for MWEI and idiom222 According to Tedeschi et al. (2022), idioms are a subset of MWEs. identification. The DiMSUM Schneider et al. (2016) dataset consists of 5,799 sentences annotated with MWEs, including, but not limited to, verbal MWEs, noun MWEs, phatics, and multi-word (MW) proper nouns. Although DiMSUM has been used in multiple MWEI studies, e.g., Kirilin et al. (2016) and Liu et al. (2021), there exist inconsistencies in the annotation, which hinders proper evaluation. Tanner and Hoffman (2023) found that over 80% of the false positives of their system were actually caused by DiMSUM’s inconsistent annotation. Next, the PARSEME corpus Savary et al. (2017); Walsh et al. (2018) is a high-quality dataset for the identification of verbal MWEs. It has been continuously updated, and the latest version (1.3) contains over 455,000 sentences across 26 languages. However, their focus is limited to verbal MWEs. Lastly, ID10M Tedeschi et al. (2022) is an idiom detection dataset consisting of automatically created training data in 10 languages and a manually curated evaluation benchmark of four languages. Their evaluation dataset was created with the help of professional annotators, but it contains only 200 sentences per language. They used Wiktionary as the source of idioms, and MWEs not in Wiktionary were skipped in the annotation. Additionally, they did not annotate discontinuous idioms.
Tagging Schemes
ID10M used the BIO scheme. DiMSUM employed the more flexible 6-tag scheme, which allows discontinuous MWEs, but cannot handle overlapping MWEs (see Section A.4 for an example of an overlapping MWE). In contrast, the parseme-tsv format Savary et al. (2017) accepts MWEs in any form, which leads us to adopt an equivalent data format.
Other Tasks
Whereas these studies addressed MWE/idiom identification as a sequence tagging task, others worked on a related but different task, idiom usage recognition. It is a binary classification of word sequences in context as idiomatic/figurative or literal. The detection task in Muzny and Zettlemoyer (2013), SemEval-2013 Task 5b Korkontzelos et al. (2013), SemEval-2022 Task 2 Subtask A Tayyar Madabushi et al. (2022), and the MAGPIE corpus Haagsma et al. (2020) are set up for this task. Our task setting is more realistic and challenging than these, requiring systems to identify all MWEs within a given sentence.
3 Task Formulation
We formulate the MWEI task as token-level sequence tagging, where each token can belong to multiple MWEs. The systems are provided with tokenized sentences. Given the sequence of tokens in each sentence, $x_{1},\ldots{},x_{n}$, the task is to output a list of MWEs where each MWE is represented as a list of word indices. The $i$-th MWE in the sentence is represented with $[\mathrm{idx}(t_{1,1}),\mathrm{idx}(t_{1,2}),\ldots{}]$ where $t_{i,j}$ is the $j$-th token of the MWE and $\mathrm{idx}(t)$ is $t$’s identifier, i.e., 1-based index. Our annotation interface allows annotations in this scheme, as illustrated in Figure 1.
4 Construction of CoAM
4.1 Data Selection and Preprocessing
In selecting the sources of our dataset, we prioritized sources aimed at general audiences so that they are in standard English and mostly free of grammatical errors. We also included both written and transcripted spoken texts. Consequently, we constructed CoAM from the following four data sources (see Section A.1 for their details). News is news text written by professional writers sourced from EMM NewsBrief, which was introduced by Glavaš and Štajner (2013).333 The dataset also contains data from WikiNews and Wikipedia, but we preferred EMM NewsBrief, which contains more MWEs according to our preliminary analysis based on the annotation by Kochmar et al. (2020). Commentary is commentaries on news from the WMT23 Shared Task monolingual training data (Kocmi et al., 2023). TED is a collection of TED talk transcriptions from two sources. (1) NAIST is the dataset by Neubig et al. (2014); (2) IWSLT is the subset of IWSLT 2017 Shared Task Cettolo et al. (2017). UD is a collection of single sentences sourced from weblogs, reviews, question-answers, newsgroups, and emails in the English Web Treebank. It is part of the Universal Dependencies (Zeman et al., 2018) and the English PARSEME corpus Walsh et al. (2018). We included the UD section of the PARSEME corpus in our dataset to enable better comparison between the annotations. We did not, however, include any sentences from UD in the test set, because UD contains user-generated content with frequent grammatical errors. All the other sources were used both for training and test splits.
For all sources, we took the first 10 (or all, when there are less than 10) sentences from the article or talk and presented them in original order. Each sentence was tokenized using SpaCy’s en_core_web_lg (3.7.1), with the exception of the UD sentences, which were already tokenized. Note that what we call words are tokens given in this manner.
4.2 Annotation Guidelines
The construction of a reliable dataset requires clear guidelines for annotators in order to minimize misannotation. However, we found no such guidelines for all-type MWE annotation, which led us to create new guidelines based on the MWE definition by Baldwin and Kim (2010) and the PARSEME annotation guidelines.444 https://parsemefr.lis-lab.fr/parseme-st-guidelines/1.3/
We define MWEs as idiomatic sequences that satisfy the following three conditions.
(a) It consists of at least two words that are always realized by the same lexemes. This condition means that his in put yourself in his shoes is not part of the MWE, as other words like Michael’s can replace his.
(b) It displays semantic, lexical, or syntactic idiomaticity. Semantic idiomaticity occurs when the meaning of an expression cannot be derived from its constituent words. It is the most important among idiomaticity types because the majority of MWEs are defined as MWEs due to this type of idiomaticity. See more description on the semantic idiomaticity in relation to non-compositionality in Section A.3. Note that transparent collocations such as stuck at are not an MWE in our definition because they are not semantically, lexically, or syntactically idiomatic. In other words, we did not take statistical idiomaticity as a qualification for an MWE, unlike Baldwin and Kim (2010), for the following reasons. First, our primary aim is to identify sequences that lead to difficulties in reading, and transparent collocations are understandable if the reader knows the constituent words. Second, statistical idiomaticity is difficult for human annotators to judge, and it would bring considerable ambiguity to the definition of MWEs.
(c) It is not a proper noun, i.e., a specific name of a person, facility, and so on. Although some previous studies include MW proper nouns in MWEs Schneider et al. (2016); Baldwin and Kim (2010), we exclude them because the identification of MW proper nouns has already been actively studied in the field of NER. Delegating MW proper noun identification to NER researchers will enable a more focused investigation on our part.
The full MWE definition is in Section A.2. The guidelines were updated whenever an issue was found, such as an ambiguous description.
Figure 1: An example sentence presented in our checkbox-based interface. Checks in the checkboxes are instantly reflected in the bordered zone at the upper right, showing which MWEs are currently marked.
4.3 Annotation Interface
Tool Flexible annota- tion Simple interface Easy collabora- tion Customiz- able interface
brat
FLAT
Ours
Table 1: Comparison of annotation tools.
For flexible and efficient annotation, we developed a novel annotation interface generator, CAIGen. CAIGen builds a checkbox-based interface in Google Sheets555 https://developers.google.com/sheets , allowing annotators to perform annotations of MWEs and other kinds of spans simply by checking checkboxes as shown in Figure 1.
Its main advantages over other comparable annotation tools are summarized in Table 1. The first tool, brat666 https://github.com/nlplab/brat Stenetorp et al. (2012), has been used for annotation of a wide range of tasks such as sentiment analysis (Pontiki et al., 2016) and NER (Tabassum et al., 2020). FLAT777 https://github.com/proycon/flat was used for the annotation of the PARSEME corpus Savary et al. (2017). The advantages of CAIGen are as follows. First, the generated interface accepts annotations of spans of any form, including discontinuous or overlapping spans (see Section A.4 for examples), making it more flexible than the previous tools. Second, CAIGen provides a simple spreadsheet-based interface, which is familiar and accessible for technical and non-technical annotators alike. Third, CAIGen alleviates researchers’ overhead of managing servers and annotator accounts by delegating to Google, while the other tools require either (1) distributing data slices to each annotator and having them run applications locally or (2) running and managing the application on a remote server. Lastly, our interface is highly customizable because CAIGen builds it using Google Apps Script, a widely used programming language based on JavaScript; for example, simple customization could allow you to collect annotations about additional information, such as span types.
4.4 Annotation
The annotation was done by (1) two annotators hired through a company and (2) five authors. The hired annotators comprise one native and one non-native English speaker, both with at least six years of experience as translators. They received feedback from the authors after annotating the first 50 or so sentences to correct misunderstandings about the annotation guidelines. We paid roughly 1 USD per sentence as a reward. Meanwhile, the author annotators consist of three native and two non-native speakers, all with a language-related degree and sufficient English proficiency to perform the task. For reliable annotation, we assigned each sentence to two annotators (one hired annotator and one author), ensuring that at least one was a native English speaker, as the task requires a rich English vocabulary.
All annotators were instructed to carefully read the guidelines and perform annotation using the checkbox-based interface. In the authors’ annotation, we marked unclear sentences to remove them later because it is hard to determine whether a span is an MWE in such sentences.
4.5 Inter-Annotator Agreement
With two annotators being assigned to each sentence, we measured the inter-annotator agreement (IAA) by the MWE-based (exact-match) F1 score. The IAA was found to be only 37.3%, which could be attributed to annotators failing to annotate valid MWEs or to subjectivity in idiomaticity judgments despite the carefully written guidelines. We thus suggest that MWE datasets should be constructed with at least two independent annotators.
4.6 Review
To solve the disagreement between the two annotators and correct any other problematic annotations, two authors—both native English speakers—reviewed all the annotations. We presented them with the annotations in a special interface for reviewing, where the tags given by only one annotator were highlighted, and we asked them to mark inappropriate tags to be deleted and newly found MWEs to be added. When the review was complete, we updated our dataset according to the marks added by reviewers.
Type PoS of Head Description Example
Noun NOUN, PRON, PROPN Noun MWEs or compounds. the middle of nowhere, red tape
Verb VERB Verbal MWEs. stand for, pick up, break a leg
Mod/ Conn ADJ, ADV, ADP, CCONJ, SCONJ Adjectival, adverbial, or connective MWEs. under the weather, of course, in spite of
Clause VERB, AUX MWEs containing (1) a verb or auxiliary verb and (2) its nominal subject, e.g., proverbs, phatic expressions. you’re welcome, the early bird gets the worm, when it comes to
Other Any MWEs whose head is not contained in them or whose PoS is none of the PoS above. and so on
Table 2: MWE types in CoAM. The PoS are denoted with UPOS tags. Underlines denote the head of an MWE.
Sentence Marked
Having never booked train tickets online before thought I would give it a try and was very surprised at how much I saved.
Would recomend giving this a a try.
Table 3: Inconsistent annotation of give_try. Note that the second sentence is reproduced as in DiMSUM, including apparent typos.
4.7 Consistency Check
One of the primary issues with the annotations in the DiMSUM dataset is their inconsistency—that is, a number of MWEs are annotated in one location and not another, despite equivalent constituents and semantics being present in both places. For example, see how give_try is labeled in only one of the sentences in Table 3.
In order to quantify this issue in both DiMSUM and CoAM—and to eliminate it from CoAM—we used a partially automated approach to find inconsistencies in both the entirety of CoAM and 1.3K sentences (the size of CoAM) randomly sampled from DiMSUM. We started by initializing an empty set $M$, then iterated through all given sentences and added all labeled MWEs from them to $M$. Next, we used a simple rule-based MWEI pipeline, repurposed from Tanner and Hoffman (2023), to find constituent groups in other sentences that could correspond to an MWE in $M$ but were not already labeled. Finally, an author and native speaker of English reviewed each of these candidate constituent groups to see if they are semantically equivalent to already labeled instances of this MWE—that is, to see if they represent an inconsistency.
We found 118 instances of inconsistencies like this in the random sample of DiMSUM and 147 in CoAM,888 When counting inconsistencies, we always consider positive labels correct, and negative labels inconsistent. reaffirming the difficulty involved in producing consistent MWE annotations. However, we then added the missing labels to all MWEs found in this way, eliminating these inconsistencies from the final CoAM data.
MWE Type Proportion (%)
Sentences Words MWEs MWE Density (%) Noun Verb Mod/Conn Clause Other
News 0360 09,328 230 5.5 33.5 47.8 16.5 0.0 2.2
Commentary 0357 09,310 272 7.0 29.8 30.5 29.4 1.1 9.2
TED 0299 06,592 212 7.2 33.0 37.3 21.2 3.8 4.7
UD 0285 05,001 160 7.2 - - - - -
Train 0780 16,817 489 6.7 33.4 35.9 22.8 1.5 6.4
Test 0521 13,414 385 6.5 30.6 40.0 22.9 1.6 4.9
Total 1,301 30,231 874 6.6 31.9 38.1 22.8 1.5 5.6
Table 4: Statistics of CoAM. MWE density is the percentage of words in MWEs. The test set comprises part of News, Commentary, and TED, while the training set comprises the rest of the data.
4.8 MWE Type Tagging
To enable fine-grained error analysis using CoAM, we automatically tagged all test-set MWEs, and training-set MWEs except those from UD with MWE types. Inspired by the classification by Schneider et al. (2014), we group MWEs into five types: Noun, Verb, Modifier/Connective (Mod/Conn), Clause, and Other. The details are described in Table 2. We tag MWEs in each sentence $s$ through the following automatic operations after dependency parsing using SpaCy. For each MWE $m$ in $s$, we look for the head word, namely, try to find $w$ that has all the other words in $m$ as its descendants. If $m$ has such $w$, we determine its type based on the PoS of $w$; for example, we tag $m$ as Noun when $w$ is a noun. If $m$ does not have such $w$, we tag $m$ as Other. We did not tag the sentences in UD because none of them are contained in the test set.
4.9 Statistics
Table 4 shows dataset statistics. CoAM has more than 1.3K sentences. The MWE density is 6–7% in both training and test sets. This proportion is much lower than that of DiMSUM Schneider et al. (2016)—13% in the grand total—most likely because they included proper noun phrases in MWEs. The right five columns show that Verb and Noun are the two most frequent MWE types across data sources. They also demonstrate that our type tagging approach successfully assigned a specific (non-Other) MWE type to almost 95% of the MWEs in CoAM.
We also report the ratio of unseen MWEs, because they are of interest to the MWE research community, with their identification being a major challenge Ramisch et al. (2020). An MWE in the test set is considered unseen if the multi-set of lemmas of its constituents was never annotated in the training set. Among 385 MWEs in our test set, 64.2% are unseen.
5 MWEI Approaches
We use CoAM to evaluate the following MWEI approaches.
5.1 MWEasWSD
MWEasWSD (MaW) is a system that uses (1) an MWE lexicon—WordNet Miller (1995)—and a rule-based pipeline to identify MWE candidates and (2) a trainable model to filter MWE candidates Tanner and Hoffman (2023). It achieved state-of-the-art performance on the DiMSUM dataset. They published four of their filtering models, among which we use the bi-encoder (BiEnc) and DCA poly-encoder (DCA). Both models are based on BERT Devlin et al. (2019), specifically bert-base-uncased. They have been trained with SemCor Miller et al. (1993), and we further fine-tune each model with the CoAM training set. We also run MaW with the rule-based pipeline only, i.e., without a filtering model.
5.2 LLM Fine-Tuning
Recent studies have shown the effectiveness of LLM fine-tuning for a wide range of tasks, such as NER Zhou et al. (2024) and grammatical error correction Kaneko and Okazaki (2023). These works achieved high performance by providing LLMs with prompts that included instructions on the task.
Inspired by their success, we perform LLM fine-tuning, where the inputs to LLMs are instructions for MWEI (a summary of the annotation guidelines) followed by formatted tokens in a sentence.
6 Experiments
6.1 Setup
Recall by MWE Type
F1 P R Noun Verb Mod/Conn Clause
(118) (154) (88) (6)
FT Llama-8B 29.4±2.1 $\mathbf{82.6}$±1.2 17.9±1.6 5.9±0.8 26.4±1.5 24.2±3.5 0.0±0.0
Llama-70B 38.4±4.8 74.5±2.6 26.1±4.6 19.2±4.2 35.7±5.7 25.4±5.6 0.0±0.0
Qwen-7B 45.2±0.6 63.2±0.9 35.2±0.8 26.0±2.0 47.2±1.0 31.8±2.0 0.0±0.0
Qwen-72B $\mathbf{55.5}$±0.5 61.5±2.6 $\mathbf{50.7}$±2.5 $\mathbf{44.4}$±3.2 $\mathbf{60.6}$±1.0 $\mathbf{50.4}$±5.1 $\mathbf{22.2}$±9.6
MaW Rule 32.7 28.3 38.7 37.3 40.9 47.7 0.0
Rule+BiEnc 41.6±0.1 48.6±0.3 36.5±0.3 32.2±0.0 41.1±0.7 44.3±0.0 0.0±0.0
Rule+DCA 42.0±0.1 48.4±0.6 37.1±0.3 33.3±0.5 40.9±0.0 45.8±0.7 0.0±0.0
Recall by Seen/Unseen Recall by In WN or Not
Seen Unseen True False
(138) (247) (163) (222)
FT Llama-8B 37.2±3.4 7.2±0.6 29.2±1.9 9.6±1.7
Llama-70B 35.0±7.3 21.1±3.2 46.8±6.5 10.8±3.2
Qwen-7B 44.4±0.8 30.0±0.8 50.1±1.4 24.2±0.5
Qwen-72B $\mathbf{58.2}$±5.8 $\mathbf{46.6}$±0.7 63.6±3.4 $\mathbf{41.3}$±1.9
MaW Rule 47.8 33.6 $\mathbf{91.4}$ 0.0
Rule+BiEnc 44.2±0.0 32.1±0.5 86.1±0.7 0.0±0.0
Rule+DCA 45.4±0.8 32.4±0.0 87.5±0.7 0.0±0.0
Table 5: Results by MWEI system and metric/MWE category, as mean percentage scores of three runs with random training seeds. The numbers in parentheses are the MWE counts. $\pm$ denotes standard deviation. The bold font denotes the highest score. Rule stands for the rule-based baseline, and WN for WordNet.
Role Message
[…]
Table 6: Example prompt for fine-tuning, based on the tsv_to_tsv format. In our main experiments, [MWE_DEFINITION] will be filled with the long MWE definition described in Section B.1.
LLM Fine-Tuning
We use four instruction-tuned LLMs that are available on Hugging Face Hub: Llama-3.1-8B-Instruct, Llama-3.1-70B-Instruct Dubey et al. (2024), Qwen-2.5-7B-Instruct, and Qwen-2.5-72B-Instruct Qwen Team (2024). We hereafter abbreviate them, e.g., to Llama-8B, omitting the versions. For efficient training and inference, we use QLoRA Dettmers et al. (2023), performing 4-bit NormalFloat quantization and double quantization. Other hyperparameters are described in Section B.1. The computational budgets are in Appendix C.
To find a performant input-output format, we conduct preliminary experiments with three formats, adjusting the prompt for each format. These experiments are detailed in Section B.1. We find tsv_to_tsv (see Table 6) is the only format the models can comply with.
All prompts contain a placeholder for a definition of MWEs to clarify what sequences should be marked. We report scores using what we refer to as the long definition in Table 5; for scores using a shorter definition, see the ablation study in Section B.1.
Evaluation Metrics
We use MWE-based precision, recall, and F1 score Savary et al. (2023). See Section B.2 for their exact definitions.
Result MWE Context Note
TP fire up The allegations have fired up the opposition, … Verb MWE in WordNet
TP at least … since at least the 1950s. Mod/Conn MWE in WordNet
FN real estate …park their toxic real estate assets … Noun MWE not in WordNet
FN you know You know, it’s very old … Clause MWE not in WordNet
Table 7: Examples of true positives (TPs) and false negatives (FNs), i.e., MWEs identified/missed in all three runs, of fine-tuned Qwen-72B.
6.2 Results and Analysis
In Table 5a, the left three columns show the overall scores. Fine-tuned Qwen-72B achieves the best F1, surpassing the highest F1 by Rule+DCA of MaW. Moreover, fine-tuned Qwen-72B outperforms all MaW systems in precision and recall. This indicates the effectiveness of fine-tuning LLMs for MWEI—particularly LLMs with a large number of parameters. One explanation for this could be that knowledge about MWEs was acquired by LLMs through pre-training, and we can harness this knowledge for MWEI by fine-tuning.
Meanwhile, all systems except for the rule-based pipeline suffer from low recall. Even the best system, fine-tuned Qwen-72B, achieves a recall of only 50.7%, failing to identify almost half of the gold MWEs. This leads us to conduct further analysis.
Analysis of Recall
The right columns in Table 5a compare recall by MWE type, showing that Clause and Noun MWEs are more difficult to identify than Mod/Conn or Verb MWEs across models. In Table 5b, the left columns show that seen MWEs are easier to identify than unseen ones across models/systems. This result resembles that of the PARSEME shared task 1.2 Ramisch et al. (2020), demonstrating that unseen MWEs remain challenging in MWEI.
The right columns in Table 5b reveal that MWEs not contained in WordNet are much more difficult to identify than those in WordNet. It is natural that MaW systems cannot identify the MWEs not in WordNet, as their rule-based pipeline cannot detect these candidates, but interestingly we find that fine-tuned LLMs also struggle to identify MWEs not in WordNet. For Qwen-72B, the F1 score gap between MWEs contained and not contained in WordNet is about 22 points, much larger than the gap between seen/unseen MWEs, about 12 points. We hypothesize that this is caused by MWEs in WordNet being more widely recognized as idioms and that this recognition is reflected in the training data of the LLMs, enhancing their ability to identify these MWEs.
Table 7 shows examples of correctly identified MWEs and missed MWEs by fine-tuned Qwen-72B. 79.0% of Verb MWEs in WordNet like fire up are correctly identified by the model. Meanwhile, the Noun MWE real estate was not identified. The aforementioned hypothesis could explain this, as real estate is not in WordNet. Accurately identifying such multi-word entities or multi-word terms Savary et al. (2019) could be a challenge for future studies.
Ablation Study on LLM Fine-Tuning
Given the high performance of fine-tuned Qwen-72B, we investigate how much fine-tuning (FT) contributes to the performance by comparing FT to zero-shot learning (ZSL) and few-shot learning (FSL). In ZSL, we provide the models with the same prompt as FT. In FSL, we sample 5 pairs of input (sentence) and output (gold MWE set) from the training set, convert them into the tsv_to_tsv format, and include them in the prompt as examples. In the sampling process, we repeat random sampling until at least two of the example sentences contain one or more MWEs so that the models have a sufficient number of examples to learn the task and format.
Table 8 shows the results. FT greatly outperforms ZSL and FSL, demonstrating the effectiveness of FT on the CoAM training set.
Llama-70B Qwen-72B
FT 38.4±4.8 55.5±0.5
FSL 4.3±0.9 14.1±0.3
ZSL 6.9 2.8
Table 8: Ablation results by MWEI method and model, as mean percentage F1 scores of three runs (for FT the randomness arises in training, and for FSL in exemplar selection). $\pm$ denotes standard deviation.
Relative Strengths of LLMs vs MWEasWSD
As shown in Table 5a, fine-tuned Qwen-72B outperforms the best MaW system, Rule+DCA, in all of F1, precision, and recall, by over 10 points. However, it does not mean fine-tuning should always be preferred. First, MaW’s encoder model (bert-base-uncased) has only 110M parameters, taking approximately $2\times 110\textrm{M}\times 4=880\textrm{M}$ bytes in total ($2$ is the number of encoders). 4-bit quantized Qwen-72B takes $72,000\textrm{M}\times 0.5=36,000\textrm{M}$ bytes, taking roughly 40 times more space than MaW. On the other hand, the performance of 4-bit quantized Qwen-7B, which takes roughly 4 times more space, is comparable with Rule+DCA, suggesting that MaW is more memory-efficient.
Furthermore, the bottleneck of MaW is identified to be the size of the lexicon Tanner and Hoffman (2023). Future work could significantly improve the system’s performance simply by using a lexicon containing more MWEs than WordNet.
7 Conclusion
In this paper, we constructed CoAM, a high-quality dataset of 1.3K sentences for MWEI covering all types of MWEs. Using a combination of human review and automated consistency checking, we addressed consistency issues that have been a problem for previous MWE datasets, enabling more accurate evaluation results for future work in MWEI. We used CoAM to evaluate two MWEI approaches: MaW and LLM fine-tuning. Our largest fine-tuned LLM performed the best, outperforming the current state-of-the-art, but all systems suffered from low recall. Consequently, we argue that MWEI overall remains a challenging task.
8 Limitations
Despite our best efforts to craft clear and unambiguous annotation guidelines, our initial annotations on CoAM suffered from very low inter-annotator agreement, suggesting possible inconsistencies. We have done our best to address this problem through the combination of human review (Section 4.6) and consistency checking (Section 4.7), but we acknowledge that in an ideal world initial inter-annotator agreement would have been higher.
Additionally, our annotation interface generator, CAIGen suffers from a few drawbacks. For example, it requires annotators to change rows when they have marked some MWEs and are marking more. Because annotators can forget to change rows, which happened in the construction of CoAM, researchers should urge them to confirm their annotations are as intended after finishing each sentence. Taking this into consideration, we release CAIGen’s source codes so that potential users can judge for themselves if its benefits outweigh its drawbacks.
9 Ethical Considerations
All sources of CoAM are permitted at minimum for use in research, as described in Section A.1. CoAM itself will be released under the condition that the users do not publish the data on the open web to prevent its leakage to the training data of future models. All annotators of CoAM consented that their annotations be made publicly accessible. Throughout the dataset construction processes involving authors, we found no contents harmful to the environment, specific groups of people, privacy, and so on.
10 Acknowledgments
To paraphrase or polish authors’ original content, we used an AI writing assistant tool, ChatGPT.999 https://chatgpt.com Financially, this work was supported by JST SPRING Grant Number JPMJSP2140.
Appendix A Construction of CoAM
A.1 Data Sources
Source Paper URL License Note
News Glavaš and Štajner (2013); Yimam et al. (2018) https://takelab.fer.hr/data/evsimplify/
https://sites.google.com/view/cwisharedtask2018/datasets
Creative Commons Attribution-NonCommercial-ShareAlike 3.0 We used the News_Train data from the CWI Shared Task 2018 datasets.
Commentary Kocmi et al. (2023) https://data.statmt.org/news-commentary/v18.1/training-monolingual/ Can be freely used for research purposes. We used the English part of v18.1.
TED (NAIST-NTT) Neubig et al. (2014) https://ahcweb01.naist.jp/old/resource/tedtreebank/ Creative Commons ShareAlike-Attribution-NonCommercial
TED (IWSLT) Cettolo et al. (2012, 2017) https://wit3.fbk.eu/2017-01 https://wit3.fbk.eu/2017-01-b Creative Commons Attribution-NonCommercial-NoDerivs 3.0
UD Bies et al. (2012); Walsh et al. (2018) https://gitlab.com/parseme/parseme_corpus_en Creative Commons ShareAlike 4.0 We used the portion of PARSEME corpus from the English Web Treebank.
Table 9: Data sources of CoAM.
A.2 MWE definition in Annotation Guidelines
In our definition, MWEs are idiomatic sequences that (a) consist of multiple words, (b) display semantic, lexical, or syntactic idiomaticity, and (c) are not proper nouns.
  1. a.
    An MWE consists of at least two words that are always realized by the same lexemes. [^10] https://en.wikipedia.org/wiki/Lexeme For example, the MWE "break up" is always realized by (1) "break" or its conjugated form such as "broke" and (2) "up". Such words cannot be replaced without distorting the meaning of the expression or violating the language conventions.
  2. b.
    An MWE displays semantic, lexical, or syntactic idiomaticity. The semantically idiomatic MWEs include "break up", the lexically idiomatic include "to and fro", and the syntactically idiomatic include "long time no see".
    • Semantic idiomaticity occurs when the meaning of an expression cannot be derived from its components. That is, you cannot necessarily infer the meaning of the expression even if you know all the senses of its components. In other words, a semantically idiomatic takes on a meaning that is unique to that combination of words.
      The inferability differs from one expression to another. The meaning of idiomatic MWEs such as "kick the bucket" cannot be inferred from its components at all. The meaning of institutionalized phrases like "traffic light" is inferable to some degree. Yet it is an MWE because it does not mean any type of light related to traffic but a specific type of light.
    • Lexical idiomaticity occurs when one or more components of an expression are not used as stand-alone words in standard English. Examples include "bide one’s time"; "bide" rarely appears by itself in today’s standard English.
    • Syntactic idiomaticity occurs when the grammatical structure of an expression cannot be derived directly from that of its components. It constitutes expressions whose grammar seems to go against standard English grammatical conventions. This includes expressions such as "all of a sudden," where "sudden" appears in its archaic noun form.
  3. c.
    An MWE is not a proper noun, such as the name of a specific person, organization, and so forth. In this project, do not annotate proper nouns unless they are part of an MWE (like Pavlov in "Pavlov’s dog"). Here is the full list of proper nouns in our definition.111111 The 11 types derive from the named entity types of OntoNotes Release 5.0 Weischedel et al. (2013).
    Type Example
    People’s names Shohei Ohtani
    Nationalities or religious or political groups African American, Sunni Muslims
    Facilities Narita International Airport, Taipei 101
    Organizations Procter & Gamble, Kyoto University
    Geopolitical entities (GPE) Los Angeles, the United Kingdom
    Non-GPE locations Mount Fuji, Amazon River
    Products Toyota Prius, Samsung Galaxy
    Named events World War II, Olympic Games Tokyo 2020
    Works of art Norwegian Wood, Bohemian Rhapsody
    Named legal documents The Magna Carta
    Named languages Middle English, American Sign Language
    Note:
    • Proper noun phrases usually start with a capital letter, while exceptions like "iPhone 15" exist.
A.3 Notes on Idiomaticity and (Non-)Compositionality
To judge whether a sequence is an MWE, we focus on (semantic) idiomaticity instead of non-compositionality, although non-compositionality has been considered an inherent property of MWEs in previous studies, such as Tedeschi et al. (2022). To discuss the difference between the two notions, consider the expression spill the beans. It is deemed compositional because it can be analyzed as being made up of spill in a “reveal” sense and the beans in a “secret(s)” sense, resulting in the overall compositional reading of “reveal the secret(s)” Sag et al. (2002). Meanwhile, the “secret(s)” sense is unique to the expression and cannot be derived from the word beans, making the whole expression semantically idiomatic. We argue that spill the beans should be identified as an MWE because the meaning of beans depends on the whole expression and that any idiomatic sequences should be identified even if they are compositional.
A.4 Annotation Interface
(a) An example of discontinuous MWE.
(b) An example of overlapping MWEs.
Figure 2: How to annotate tricky MWEs using our interface.
Our interface generated by CAIGen enables annotators to mark spans in any form, including discontinuous and overlapping ones, only by checking checkboxes. Figure 2 shows the examples.
Appendix B Experimental Setup
B.1 LLM fine-tuning
LoRA $r$ 64
$\alpha$ 16
Dropout 0.05
Target modules All linear layers
Training Epoch 3
Effective batch size 32
Learning rate 2e-4
Learning rate scheduler constant
Optimizer paged_adamw_8bit ($\beta_{2}$ = 0.999)
Max grad norm 0.3
Table 10: Hyperparameters for fine-tuning. The epoch was determined based on preliminary experiments, while other parameters are based on Dettmers et al. (2023).
Name Example (Input $\rightarrow$ Output)
[…]
str_to_str_ number_span ACL stands for Association for Computational Linguists . $\rightarrow$ ACL <1>stands for</1> Association for Computational Linguists .
[…]
Table 11: Input-output formats for fine-tuning. Bold fonts denote the format employed for the main experiments. Suppose the sentence given as the example is ACL stands for Association for Computational Linguistics.
Hyperparameters
Table 10 shows hyperparameters used for training and evaluation. At inference, we perform greedy decoding.
Input-Output Format
To investigate the optimal input-output format, we perform preliminary experiments. We prepare three input-output formats described in Table 11. All the output formats allow us to represent MWEs in any form, including discontinuous or overlapping MWEs. We train and evaluate Llama-8B and Qwen-7B with CoAM, using each input-output format. We use the hyperparameters described in Table 10 and the long MWE definition (same as the main experiments).
As a result, tsv_to_tsv turned out to be the only format the models can comply with. Employing other formats results in the violation of the format despite carefully written instructions. With dict_to_dict_list, the models produce wrong indices in the outputs. With str_to_str_number_span, the models delete spaces before punctuations.
Name Content
[…]
Table 12: MWE definitions to be included in prompts for fine-tuning.
R by MWE Type
F1 P R Noun Verb Mod/Conn Clause
(118) (154) (88) (6)
Long Llama-8B 29.4±2.1 $82.6$±1.2 17.9±1.6 5.9±0.8 26.4±1.5 24.2±3.5 0.0±0.0
Qwen-7B 45.2±0.6 63.2±0.9 35.2±0.8 26.0±2.0 47.2±1.0 31.8±2.0 0.0±0.0
Short Llama-8B 8.6±14.8 28.9±50.0 5.0±8.7 1.4±2.4 7.1±12.4 7.2±12.5 $\mathbf{0.0}$±0.0
Qwen-7B 32.8±0.2 $70.8$±1.6 21.4±0.1 12.1±0.5 30.5±0.6 22.3±0.7 0.0±0.0
Table 13: Ablation results, as mean percentage scores of three runs with random seeds. The numbers in parentheses are the MWE counts. $\pm$ denotes standard deviation.
MWE Definition
To validate the efficacy of the long MWE definition, we perform an ablation study comparing the long definition to the short definition. They are both a summary of the full definition described in Section A.2. As shown in Table 12, the long definition has 162 words while the short is further contracted to 57 words. We perform experiments with Llama-8B and Qwen-7B, using the tsv_to_tsv format and the hyperparameters described in Table 10 (same as the main experiments).
Table 13 presents the result. For both models, the long definition achieves higher F1 scores than the short by more than 12 points, indicating the effectiveness of a detailed definition.
B.2 Evaluation Metrics
We use MWE-based (exact-match) precision, recall, and F1 score Savary et al. (2023). Let $G$ be the set of gold MWEs and $H$ the set of predicted MWEs (hypothesis),
$\mathrm{Recall}$ $=|G\cap H|/|G|$ (1)
$\mathrm{Precision}$ $=|G\cap H|/|H|$ (2)
where each MWE is represented with the ID of the sentence and the IDs of its constituent tokens. F1 is the harmonic mean of precision and recall.
Appendix C Computational Budgets
In experiments for MaW (Rule+BiEnc and Rule+DCA), we use a single NVIDIA RTX 2080Ti GPU with 11GB RAM. Each run of training and testing takes around 8 minutes. Consequently, the total GPU hours for MaW are estimated to be 0.8 hours.
In LLM fine-tuning, we use NVIDIA RTX 6000 GPU with 48GB RAM for smaller models (Llama-8B and Qwen-7B) and NVIDIA A100 PCIe with 80GB RAM for larger models (Llama-70B and Qwen-72B). Each run of training and testing takes around 66 minutes for smaller models and 588 minutes for larger models. Thus, the total GPU hours for fine-tuning experiments of smaller and larger models are estimated to be 6.6 hours and 58.8 hours, respectively.